# Cross-composition Feature Disentanglement for Compositional Zero-shot Learning

### v1

@@ Introduction
Specifically, as shown in Figure 2(b), given a target image $x_{(a,o)}$ to predict, which is labeled by $c=(a,o)$, we first randomly sample an auxiliary composition that shares the same attribute as $x_{(a,o)}$ but has different object $o^{\prime}$, and select one of its images $x_{(a,o^{\prime})}$ as an auxiliary image. Then, we feed these two images into CLIP to output two hidden states $\bm{H}_{(a,o)}^{v}$ and $\bm{H}_{(a,o^{\prime})}^{v}$ and compute the cross-attention as, with $\bm{H}_{(a,o^{\prime})}^{v}$ as the query and $\bm{H}_{(a,o)}^{v}$ as the key and value:
$CrossAttention(\bm{Q,K,V})=softmax(\tfrac{\bm{Q}\bm{K}^{T}}{\sqrt{d^{\prime}}})\bm{V}$ (4)
@@
Notably, the above cross-attention can only process two attribute-sharing compositions at one time. To introduce more compositions to learn more general attribute features, the model relies on the batched data and random sampling to switch the auxiliary composition. However, when an attribute is diverse with extensive composition neighbors, e.g., broken, which also means extensive candidate auxiliary compositions, the model requires more switches to traverse them, leading to inferior overall performance as shown in Table 1. To balance the switch times of attributes with different numbers of neighbors, we propose to select some representative compositions instead of all the compositions as the candidates, and for a target image $x_{(a,o)}$, the top-$n$ objects that are most and least relevant to the target object $o$ are paired with the target attribute $a$ to serve as the representative (auxiliary) compositions, while the relevance between two objects can be determined by the number of common attributes that co-occur in their associated compositions in the training set; for example, if there are compositions red tomato and red apple in the training set, red is a common attribute of the objects tomato and apple.
@@
To this end, we refer to the attribute-object edges in the training compositional graph to first create an att-obj graph with structure matrix $A^{\text{att-obj}}\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{O}|}$, where $A^{\text{att-obj}}_{i,j}=1$ means attribute $i$ and object $j$ form a valid seen composition while 00 not. Then, an object relevance graph can be created and its structure matrix is found as $A^{\text{obj}}=(A^{\text{att-obj}})^{T}A^{\text{att-obj}}$ with size $|\mathcal{O}|\times|\mathcal{O}|$, where $A^{\text{obj}}_{i,j}$ represents the number of common attributes between $i$-th and $j$-th objects, large numbers mean higher relevance. With this relevance graph, for target image $x_{(a,o)}$, and all of its $a$-sharing compositions besides $(a,o)$, denoted as $\{(a,o^{\prime})\}$, we next refer to $A^{\text{obj}}_{o}$ to obtain the relevance scores of objects in $\{o^{\prime}\}$, and select the compositions whose objects have top-$n$ maximum and top-$n$ minimum non-zero scores as the representative compositions. Figure 2(b) presents a running example of this procedure. We perform a weighted random sampling over these representative compositions, i.e., the probability for selecting $(a,o^{\prime})$ is determined by the normalized relevance score between $o^{\prime}$ and $o$, more details are attached in Appendix A.
@@
The same applies to the output of the object cross-attention $\bm{H}_{(a,o)\rightarrow O}^{v}$, which is learned from a set of representative $o$-sharing compositions selected by referring to the attribute relevance matrix $A^{\text{att}}=A^{\text{att-obj}}(A^{\text{att-obj}})^{T}$ and its row value $A^{\text{att}}_{a}$. In this way, we learn the cross-composition-sharing primitive features of an input image, and have the updated image features: $\bm{\widetilde{H}}_{(a,o)}^{v}=\bm{H}_{(a,o)\rightarrow A}^{v}+\bm{H}_{(a,o)\rightarrow O}^{v}$, which will be taken as the final output of one V-Adapter together with $\bm{H}_{(a,o)\rightarrow A}^{v}$ and $\bm{H}_{(a,o)\rightarrow O}^{v}$.
Integrating Adapters into CLIP
@@
Given a single L-Adapter and V-Adapter, we next present how to insert them into two connected computation units in CLIP. Inspired by ViT (Dosovitskiy et al. 2020) which tends to learn general features at lower layers and learn specific features at higher layers, we add our adapters starting from the top transformer blocks. After preliminary validations on MIT-States (see our Ablation Study results for more), we decided to i) insert adapters at the last three transformer blocks of both text and image encoders, and ii) add L-Adapters behind the self-attention layer and feed-forward layer in each language transformer block (i.e., the positions ③ and ② in Figure 2(c)), and attach V-Adapters after the whole vision transformer block (i.e., the position ① in Figure 2(c)). Moreover, we build a skip connection between the input and the output of each adapter before inputting it into the next computation unit of CLIP. We use CLIP’s pre-trained word embeddings to initialize each word in our prompts, including the prefix words “a photo of”, and keep these word embeddings trainable for capturing more task-specific knowledge.
Training and Inference
@@
$s(x_{i},c_{i}=(a_{i},o_{i}))=\alpha\ [\bm{\widehat{h}}^{v}_{i}\cdot\bm{\widehat{h}}^{t}_{c_{i}}]+\beta\ [\bm{\widehat{h}}^{v}_{i\rightarrow A}\cdot\bm{\widehat{h}}^{t}_{a_{i}}]$ (6)
@@
We optimize these adapters and trainable token embeddings by minimizing the cross-entropy loss on the training set $\mathcal{D}_{tr}$ with seen compositions from $\mathcal{C}_{s}$:
@@
where $\tau$ is the temperature commonly used in CLIP, $c_{i}=(a_{i},o_{i})$ is expected to be the ground-truth label.
During inference, for each testing image $x_{t}$, we estimate the compatibility score between $x_{t}$ and each testing composition $c=(a,o)$ from $\mathcal{C}_{te}$ as Equation 6. The composition that has the highest compatibility score is the predicted label. Besides, since we have no idea about the attribute and object labeled for the testing image, we take itself as the primitive-sharing images for computing $\bm{H}_{t\rightarrow A}^{v}$ and $\bm{H}_{t\rightarrow O}^{v}$.
@@
Experiments
Experimental Settings
Datasets and Metrics
@@
We experiment with three popular benchmarks for CZSL: MIT-States (Isola, Lim, and Adelson 2015), UT-Zappos (Yu and Grauman 2014) and C-GQA (Naeem et al. 2021), and follow (Purushwalkam et al. 2019; Naeem et al. 2021) to split the data for training, validation and testing. For each dataset, we compute the prediction accuracy of seen and unseen compositions, and report four metrics: the best Seen (S) and Unseen (U) accuracy, the best harmonic mean (H), and the Area Under the accuracy Curve (AUC). Among them, AUC is the most comprehensive one and is widely adopted as the core metric by previous works (Purushwalkam et al. 2019; Zheng, Zhu, and Nevatia 2024). Please see Appendix B for more datasets and metrics details.
Closed World and Open World Settings
@@
Given the attribute set $\mathcal{A}$ and object set $\mathcal{O}$, the complete compositional label set $\mathcal{C}$ should be the Cartesian product of $\mathcal{A}$ and $\mathcal{O}$, i.e., $\mathcal{C}=\mathcal{A}\times\mathcal{O}$ with size $|\mathcal{A}|\times|\mathcal{O}|$. However, current benchmarks often operate in a closed world setting where unseen compositions $\mathcal{C}_{u}$ in $\mathcal{C}_{te}$ are a small subset of $\mathcal{C}\setminus\mathcal{C}_{s}$ and are assumed to be known. For example, MIT-States contains 28,175 possible compositions with 115 attributes and 245 objects, but the label space for testing is limited to 1,962 compositions (1,262 seen and 700 unseen), covering less than 7% of the complete set. Thus, we follow (Mancini et al. 2021) to evaluate our model in the open world setting, where the testing images remain unchanged but the testing label space is all possible combinations, i.e., $\mathcal{C}_{te}=\mathcal{C}$ and $\mathcal{C}_{u}=\mathcal{C}\setminus\mathcal{C}_{s}$, which is more challenging as the models have to generalize from a small set of seen to a very large set of unseen compositions. Notably, not all the combinations are feasible, such as eroded cat, for this, we apply post-training calibration (Nayak, Yu, and Bach 2023; Xu, Chai, and Kordjamshidi 2024) to filter out unreasonable compositions.
Baselines and Model Variants
@@
We mainly compare our DCDA with the existing CLIP-based CZSL methods, including the vanilla CLIP without fine-tuning, CSP (Nayak, Yu, and Bach 2023), HPL (Wang et al. 2023a), DFSP (Lu et al. 2023), GIPCOL (Xu, Chai, and Kordjamshidi 2024), CAILA (Zheng, Zhu, and Nevatia 2024), and Troika (Huang et al. 2024). We also include two non-CLIP-based baselines that are most similar to us, namely CGE (Naeem et al. 2021) and ADE (Hao, Han, and Wong 2023).
@@
In V-Adapters, we propose a novel primitive-relevance guided (PRG) sampling method to select representative auxiliary compositions to sample rather than performing purely random (RD) sampling over all neighboring compositions. For detailed comparisons, we develop a variant DCDA[RD] and denote the vanilla model as DCDA[PRG]. Given the imbalanced sample distribution, we also focus on the tail compositions in the neighbor set, and sample them according to the reciprocal of their image numbers (N) to supply DCDA[PRG], leading to a new variant DCDA[PRG+N], these two sampling strategies are switched batch by batch.
Implementation Details
We implement our models with PyTorch and use Adam as the optimizer, with the learning rate set to 5e-5, 5e-5, 1e-5, and the batch size set to 32, 32, 16 for MIT-States, UT-Zappos, C-GQA, respectively, the weight decay is set to 5e-5 for all datasets. The GNN module is implemented as GCN with $K=2$. $n$ for selecting the representative auxiliary compositions is set to 5. The initialized value of $\alpha,\beta,\gamma$ are all set to 1, and later optimized together with other parameters. The optimum configurations w.r.t these hyper-parameters are determined by AUC on the validation set. All the experiments are run on a single NVIDIA Tesla A100 GPU with 40GB memory.
Main Results
The overall prediction results in closed world are reported in the top section of Table 1. It can be seen that our models DCDA[PRG] and DCDA[PRG+N] always outperform the baselines of both non-CLIP-based and CLIP-based. Especially, on MIT-States, DCDA[PRG] and DCDA[PRG+N] achieved the top-2 best results and have been the new SOTA, on all metrics. They outperform the previous SOTA CAILA by a large margin, with over 3% improvements on both AUC and H, and over 6% improvements on seen accuracy.
On UT-Zappos, DCDA[PRG+N] is the best with 3 out of 4 metrics. Despite the inferior results on H compared with CGE, it beats CGE a lot on the core metric AUC. In contrast to the slight performance gap between DCDA[PRG+N] and DCDA[PRG] on MIT-States, DCDA[PRG+N] shows great superiority on UT-Zappos, which is consistent with the fact that UT-Zappos is a more imbalanced dataset (the standard deviation of its sample size for all classes is $\sim$465, which is larger than that of MIT-States ($\sim$12) and C-GQA ($\sim$22)).
In our preliminary experiments, we found our models underperform the previous SOTA CAILA on C-GQA. This may be because the images in C-GQA have a lower resolution and smaller sizes, which might require more task-specific parameters to fit, just as CAILA inserts adapters into all the transformer layers in CLIP, while we only insert into the last three layers. To this end, after multiple attempts, we finally decided to freeze the image encoder to fully update its parameters together with our V-Adapters (please see Table 5 in Appendix B for more). The resulting DCDA[PRG] and DCDA[PRG+N] eventually achieved very competitive results against CAILA, as Table 1 shows. Here, to afford more trainable parameters, we run our models and some baselines with CLIP-base on C-GQA, while on MIT-States and UT-Zappos, we still follow previous works to use CLIP-large.
Compared with DCDA[RD], DCDA[PRG] achieves great outperformance on all datasets, indicating the effectiveness of our proposed primitive-relevance guided sampling strategies. Especially, the improvement on MIT-States is promising, since the number of compositions surrounding its each primitive is more imbalanced than that of UT-Zappos. Although C-GQA is also more imbalanced, the improvement is limited by its low-quality and small-size image samples.
Among all CLIP-based methods, CAILA, Troika and our DCDA are the only three that propose to insert trainable adapters into CLIP’s image encoder to learn disentangled image features, and achieve the top-3 best results on all datasets with significant improvements over other methods. This indicates the disentanglement of visual primitive features plays a key role in CZSL, which also facilitates more effective solutions to adapt CLIP to the CZSL task well, such as Troika’s multi-path solution. Also, our method consistently beats CAILA and Troika on both the general dataset MIT-States and the domain-specific dataset UT-Zappos. The comparable results on C-GQA also motivate us to develop more advanced learning paradigms in the future.
Table 1’s bottom section records the results in open world, from which we have similar observations as in closed world: our method achieves new SOTA on MIT-States and UT-Zappos, and is still the second best on C-GQA, illustrating that our method has a good generalization ability even though the unseen label space is extremely enlarged.
Effectiveness of Adapters
The whole Adapters
We evaluate the contribution of L-Adapter and V-Adapter by analyzing the performance drop when one of them is removed. Notably, when all L-Adapters (reps. V-Adapters) are removed, the text (resp. image) encoder will act like CLIP’s default frozen encoders to output an entangled representation for each input prompt (resp. image). We conduct experiments on MIT-States dataset under the closed world setting, the results are shown in the second and third lines of Table 2. We can see that the performance both declines when L-Adapters or V-Adapters are removed, indicating that they two both have a positive contribution to the DCDA model and are complementary to each other. We also observe that the performance decrease of removing V-Adapters is greater than that of removing L-Adapters, which is consistent with our statement: textual primitive features are less entangled than visual ones, and the independent textual primitive features can still be captured by setting individual primitive prompts.

### v2

@@ 3.2 The design of V-Adapter
Specifically, as shown in Figure 3(b), given a target image $x_{(a,o)}$ to predict, which is labeled by $c=(a,o)$, we first randomly sample an auxiliary composition that shares the same attribute as $x_{(a,o)}$ but has different object $o^{\prime}$, and select one of its images $x_{(a,o^{\prime})}$ as an auxiliary image [^2] Regarding the absence of attribute-sharing compositions or images, let’s consider an input image of red apple, if there are no other red objects in the dataset, we take red apple itself as the auxiliary composition and sample one of its samples (excluding the input one) as the auxiliary image. And, if there are no other red apple images, we treat the input image itself as the auxiliary image.. Then, we feed these two images into CLIP to output two hidden states $\bm{H}_{(a,o)}^{v}$ and $\bm{H}_{(a,o^{\prime})}^{v}$ and compute the cross-attention as, with $\bm{H}_{(a,o^{\prime})}^{v}$ as the query and $\bm{H}_{(a,o)}^{v}$ as the key and value:
$\text{CrossAttention}(\bm{Q,K,V})=softmax(\tfrac{\bm{Q}\bm{K}^{T}}{\sqrt{d^{\prime}}})\bm{V}$ (4)
@@
Notably, the above cross-attention can only process two attribute-sharing compositions at one time. To introduce more compositions to learn more general attribute features, the model relies on the batched data and random sampling to switch the auxiliary composition. However, when an attribute is diverse with extensive composition neighbors, e.g., red or broken, which also means extensive candidate auxiliary compositions, the model requires more switches to traverse them, leading to inferior overall performance as shown in Table 1. To balance the switch times of attributes with different numbers of neighbors, we propose to select some representative compositions instead of all the compositions as the candidates, and for a target image $x_{(a,o)}$, the top-$n$ objects that are most and least relevant to the target object $o$ are paired with the target attribute $a$ to serve as the representative auxiliary compositions, while the relevance between two objects can be determined by the number of common attributes that co-occur in their associated compositions in the training set; for example, if there are compositions red tomato and red apple in the training set, red is a common attribute of the objects tomato and apple.
@@
To this end, we refer to the attribute-object edges in the training compositional graph to first create an att-obj graph with structure matrix $A^{\text{att-obj}}\in\mathbb{R}^{|\mathcal{A}|\times|\mathcal{O}|}$, where $A^{\text{att-obj}}_{i,j}=1$ means attribute $i$ and object $j$ form a valid seen composition while 00 not. Then, an object relevance graph can be created and its structure matrix is found as $A^{\text{obj}}=(A^{\text{att-obj}})^{T}A^{\text{att-obj}}$ with size $|\mathcal{O}|\times|\mathcal{O}|$, where $A^{\text{obj}}_{i,j}$ represents the number of common attributes between $i$-th and $j$-th objects, large numbers mean higher relevance. With this relevance graph, for target image $x_{(a,o)}$, and all of its $a$-sharing compositions besides $(a,o)$, denoted as $\{(a,o^{\prime})\}$, we next refer to $A^{\text{obj}}_{o}$ to obtain the relevance scores of objects in $\{o^{\prime}\}$ w.r.t $o$, and select the compositions whose objects have top-$n$ maximum and top-$n$ minimum non-zero scores as the representative compositions. Figure 3(b) presents a running example of this procedure. Finally, we perform a weighted random sampling over these representative compositions, i.e., the probability for selecting $(a,o^{\prime})$ is determined by the normalized relevance score between $o^{\prime}$ and $o$, more details are attached in Appendix A.1.
@@
The same applies to the output of the object cross-attention $\bm{H}_{(a,o)\rightarrow O}^{v}$, which is learned from a set of representative $o$-sharing compositions selected by referring to the attribute relevance matrix $A^{\text{att}}=A^{\text{att-obj}}(A^{\text{att-obj}})^{T}$ and its row value $A^{\text{att}}_{a}$. In this way, we learn the cross-composition-sharing primitive features of an input image, and have the updated image features: $\bm{\widetilde{H}}_{(a,o)}^{v}=\bm{H}_{(a,o)\rightarrow A}^{v}+\bm{H}_{(a,o)\rightarrow O}^{v}$, which will be the final output of one V-Adapter together with $\bm{H}_{(a,o)\rightarrow A}^{v}$ and $\bm{H}_{(a,o)\rightarrow O}^{v}$.
3.3 Integrating Adapters into CLIP
@@
Given a single L-Adapter and V-Adapter, we next present how to insert them into two connected computation units in CLIP. Inspired by ViT Dosovitskiy et al. (2020) which tends to learn general features at lower layers and learn specific features at higher layers, we add our adapters starting from the top transformer blocks. After preliminary validations on MIT-States (see our Ablation Study results for more), we decided to i) insert adapters at the last three transformer blocks of both text and image encoders, and ii) add L-Adapters behind the self-attention layer and feed-forward layer in each language transformer block (i.e., the positions ③ and ② in Figure 3(c)), and attach V-Adapters after the whole vision transformer block (i.e., the position ① in Figure 3(c)). Moreover, we build a skip connection between the input and the output of each adapter before inputting it into the next computation unit of CLIP. We use CLIP’s pre-trained word embeddings to initialize each word in our prompts, including the prefix words “a photo of”, and keep these word embeddings trainable for capturing more task-specific knowledge.
3.4 Training and Inference
@@
$s(x_{i},c_{i}=(a_{i},o_{i}))$ $=\alpha\ [\bm{\widehat{h}}^{v}_{i}\cdot\bm{\widehat{h}}^{t}_{c_{i}}]$ (6)
$+\beta\ [\bm{\widehat{h}}^{v}_{i\rightarrow A}\cdot\bm{\widehat{h}}^{t}_{a_{i}}]$
@@
We optimize these adapters and trainable token embeddings by minimizing the cross-entropy loss on the training set $\mathcal{D}_{tr}$ with seen compositions from $\mathcal{C}_{s}$, with $\tau$ as the temperature widely used in CLIP, and $c_{i}=(a_{i},o_{i})$ as the ground-truth label:
@@
During inference, for each testing image $x_{t}$, we estimate the compatibility score between $x_{t}$ and each testing composition $c=(a,o)$ from $\mathcal{C}_{te}$ as Equation 6. The composition that has the highest compatibility score is the predicted label. Besides, since we have no idea about the attribute and object labeled for $x_{t}$, we take itself as the primitive-sharing images to compute $\bm{H}_{t\rightarrow A}^{v}$ and $\bm{H}_{t\rightarrow O}^{v}$.
@@
4 Experiments
4.1 Experimental Settings
@@
Datasets and Metrics. We experiment with three popular benchmarks for CZSL: MIT-States Isola et al. (2015), UT-Zappos Yu and Grauman (2014) and C-GQA Naeem et al. (2021), and follow Purushwalkam et al. (2019); Naeem et al. (2021) to split the data for training, validation and testing. For each dataset, we compute the prediction accuracy of seen and unseen compositions, and report four metrics: the best Seen (S) and Unseen (U) accuracy, the best harmonic mean (H), and the Area Under the accuracy Curve (AUC). Among them, AUC is the most comprehensive one and is widely adopted as the core metric by previous works Purushwalkam et al. (2019); Zheng et al. (2024). Please see Appendix B for more datasets and metrics details.
@@
Closed World and Open World Settings. Given the attribute set $\mathcal{A}$ and object set $\mathcal{O}$, the complete compositional label set $\mathcal{C}$ should be the Cartesian product of $\mathcal{A}$ and $\mathcal{O}$, i.e., $\mathcal{C}=\mathcal{A}\times\mathcal{O}$ with size $|\mathcal{A}|\times|\mathcal{O}|$. However, current benchmarks often operate in a closed world setting where unseen compositions $\mathcal{C}_{u}$ in $\mathcal{C}_{te}$ are a small subset of $\mathcal{C}\setminus\mathcal{C}_{s}$ and are assumed to be known. For example, MIT-States contains 28,175 possible compositions with 115 attributes and 245 objects, but the label space for testing is limited to 1,962 compositions (1,262 seen and 700 unseen), covering less than 7% of the complete set. Thus, we follow Mancini et al. (2021) to evaluate our model in the open world setting, where the testing images remain unchanged but the testing label space is all possible combinations, i.e., $\mathcal{C}_{te}=\mathcal{C}$ and $\mathcal{C}_{u}=\mathcal{C}\setminus\mathcal{C}_{s}$, which is more challenging as the models have to generalize from a small set of seen to a very large set of unseen compositions. Notably, not all the combinations are feasible, such as eroded cat, for this, we apply post-training calibration Nayak et al. (2023); Xu et al. (2024) to filter out unreasonable compositions.
@@
Baselines and Model Variants. We mainly compare our DCDA with the existing CLIP-based CZSL methods, including the vanilla CLIP without fine-tuning, CSP Nayak et al. (2023), HPL Wang et al. (2023a), DFSP Lu et al. (2023), CAILA Zheng et al. (2024), and Troika Huang et al. (2024). We also include two non-CLIP-based baselines that are most similar to us, namely CGE Naeem et al. (2021) and ADE Hao et al. (2023).
@@
In V-Adapters, we propose a novel primitive-relevance guided (PRG) sampling method to select representative auxiliary compositions to sample rather than performing purely random (RD) sampling over all neighboring compositions. For detailed comparisons, we develop a variant DCDA[RD] and denote the vanilla model as DCDA[PRG]. With the imbalanced sample distribution, we also focus on the tail compositions in the neighbor set, and sample them according to the reciprocal of their image numbers (N) to supply DCDA[PRG], leading to a new variant DCDA[PRG+N], these two sampling strategies are switched batch by batch. Please see Appendix B.3 for more implementation details.
4.2 Main Results
Closed-world Performance. As shown in Table 1 (top), our DCDA variants achieve state-of-the-art performance across benchmarks. On MIT-States, both DCDA[PRG] and DCDA[PRG+N] surpass previous methods with over 3% improvements in AUC and harmonic mean (H), and over 6% gains in seen accuracy compared to CAILA. For UT-Zappos, DCDA[PRG+N] is the best on three metrics and achieves substantial AUC improvements over CGE despite the inferior result on H. The initial experiments showed suboptimal performance on C-GQA due to its low-quality images, unfreezing CLIP’s image encoder eventually enabled DCDA to achieve competitive second-place results, see Appendix B.4 for more.
Model Variants Analyses. The significant margin between DCDA[PRG] and DCDA[RD] on all datasets validates our primitive-relevance guided sampling strategy. Especially, the improvement on MIT-States is promising, since the number of compositions surrounding its each primitive is more imbalanced than that of UT-Zappos. Moreover, in contrast to the slight performance gap between DCDA[PRG+N] and DCDA[PRG] on MIT-States and C-GQA, DCDA[PRG+N] shows great superiority on UT-Zappos, demonstrating superior handling of class imbalance (sample size std: UT-Zappos’s 465 vs. MIT-States’ 12 vs. C-GQA’s 22).
Cross-method Insights. Among CLIP-based methods, CAILA, Troika, and DCDA form the top tier by injecting adapters for visual disentanglement. Our approach outperforms both competitors on both the general dataset MIT-States and the domain-specific dataset UT-Zappos with fewer trainable parameters through partial-layer adapter insertion (our last-3 layers vs. full layers in CAILA and Troika), demonstrating superior cross-domain adaptability, and preserving CLIP’s generalization as evidenced by open-world results (Table 1, bottom). The comparable results on C-GQA also motivate us to develop more advanced learning paradigms in the future.